OpenStudy (kaederfds):

y=sin( x^2 y^2 ) y'=?

OpenStudy (jiteshmeghwal9):

apply chain rule

OpenStudy (mhchen):

Any ideas?

OpenStudy (kaederfds):

@jiteshmeghwal9 could you please explain more?

OpenStudy (jiteshmeghwal9):

assume x^2y^2=u sin(x^2y^2)=sinu \[\frac{d (sinu)}{du} * \frac{du}{dx}\]

OpenStudy (jiteshmeghwal9):

\[\frac{dy}{dx}=\frac{d(sinu)}{du}*\frac{d(x^2y^2)}{dx}\]